OpenStudy (anonymous):

An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the objects maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t² + 64t + 80? Round to the nearest foot.

OpenStudy (anonymous):

max is at vertex. first coordinate of the vertex is \(-\frac{b}{2a}\) in this case it is \(-\frac{64}{2\times -16}=2\)

OpenStudy (anonymous):

replace \(t\) by 2 to get your answer

OpenStudy (anonymous):

144ft?